Two conducting spheres of radii R1 and R2 are kept widely separated from each other. What are their individual capacitances? If the spheres are connected by a metal wire, what will be the capacitance of the combination? Think in terms of series-parallel connections.

Grade:11

1 Answers

Kevin Nash
9 years ago
Sol. C = q/V , Now V = Kq/R So, C base 1 = q/(Kq/R base 1) = R base 1/K = 4 πε base 0R base 1 Similarly c base 2 = 4 πε base 0R base 2 The combination is necessarily parallel. Hence Ceq = 4 πε base 0R base 1 + 4 πε base 0R base 2 = 4 πε base 0(R base 1 + R base 2)